OpenStudy (anonymous):

what is 1.3 moles of calcium converted to grams?

OpenStudy (jfraser):

what's the molar mass of calcium? (from the periodic table)

OpenStudy (chmvijay):

Grams = moles *molar mass

OpenStudy (anonymous):

52.1grams. moles = mass (in grams)/RMM (relative molecular mass) 1.3 moles = mass (in grams)/ 40.078 mass = 40.078 x 1.3 = 52.1grams